OpenStudy (ellamoyseyuk):

A real estate company sells plots of land. The plot shown at the right costs $84,120. What is the price per square foot of the land? Explain how you found your answer.

OpenStudy (aaronandyson):

Any other information given?

OpenStudy (ellamoyseyuk):

yeah sorry about that, i just posted it.

OpenStudy (triciaal):

first find the area of the plot of land. the area has units in sq ft. let's say the area of the plot is P given the cost is $84,120 for P sq ft the question is asking the price for 1 sq ft

OpenStudy (phi):

this looks a bit painful. One way is to find the area of the rectangle around the plot then subtract off the "empty" parts if we can figure them out.

OpenStudy (phi):

Though looking at it some more, maybe two trapezoids is a lot easier

OpenStudy (ellamoyseyuk):



OpenStudy (ellamoyseyuk):

i don't get how to get the other two side @phi

OpenStudy (phi):

the area of a trapezoid is the average of the top and bottom times the height. you have the top and bottom: we add them up and divide by 2 now we need the height. (which we can figure out... do you see how ?)

OpenStudy (phi):

the total height of the figure is 260, and the lower shape is 150 tall. that means the rest is the height of the top figure.

OpenStudy (phi):

for the bottom trapezoid, we have to know that the "little mark" on the right side and the bottom means those two sides are the same length. so we have top = 265, bottom is 150, and height is 150 the area is ½ (top + bottom) * height

OpenStudy (ellamoyseyuk):

Thre first rapiozd is A = 1/2(265+125)(150) a = 1/2(390)(150) a =29250 A = 1/2(265+150)(150) = A = 1/2(415)(150) = A = 31125 A =29250+31125= 60375 ft^2

OpenStudy (ellamoyseyuk):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The height of the top trapezoid is not 150

OpenStudy (ellamoyseyuk):

is it 130

jimthompson5910 (jim_thompson5910):

260 - 150 = ??

OpenStudy (ellamoyseyuk):

260 - 150 =110

jimthompson5910 (jim_thompson5910):

yes 110 is the height of the top trapezoid

OpenStudy (ellamoyseyuk):

A = 1/2(265+110)(150) a = 1/2(375)(150) a =28125 A = 1/2(265+150)(150) = A = 1/2(415)(150) = A = 31125 A =28125+31125= 59250 ft^2

jimthompson5910 (jim_thompson5910):

Area of top trapezoid A = (1/2)*(b1+b2)*h A = (1/2)*(265+125)*110 A = ??

OpenStudy (ellamoyseyuk):

A = 1/2(265+125)(110) a = 1/2(390)(110) a =21450 A = 1/2(265+150)(150) = A = 1/2(415)(150) = A = 31125 A =21450+31125= 52575 ft^2

jimthompson5910 (jim_thompson5910):

52575 is the total area of the figure

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

price per square foot = (total cost)/(total area)

OpenStudy (ellamoyseyuk):

84120 divide it by 52575 =1.6

jimthompson5910 (jim_thompson5910):

So the cost per square foot is $1.60 Each square foot of land costs $1.60